cancelled 9-2 tonight : Beam current

Dong-Woon Shin dwshin at stanford.edu
Wed Feb 11 13:57:28 PST 2004


Hi.

I also found beam current fluctuating when I was measuring the faraday cup last 
night.  I think it moved from 0.68-1.02 nA af first.  (usu. 1.42)  It became stablized after 
couple of measurements.  I moved the stage to other positions, came back to the 
faraday cup and remeasured.  Beam current was staying the same.  However the 
value stayed low.  I hope it can be resolved after the bake as James suggested.

- Dongwoon


인용 Lindsay Moore <lsmoore at stanford.edu>:

> also, whatever list raith-pcs at snf.stanford.edu is, it does not send email
> 
> to the whole raith list.
> 
> 
> 
>   here is the info from charis:
> 1) Beam current low and fluctuating *a lot*. Last user (John Paul)
> records 
> 30pm at 30 micron aperture (compared to 140 usuall). I got 63 +/- 5. At
> 120 
> micron aperture,current was 1.252nA and decreased by 0.01 nA over a five
> 
> minute period. I noticed from the logbook that the current has been 
> steadily decreasing over the past 2 days. 2) Column not moving. After 
> write, joystick did not move column. Nor did software. I managed to 'fix'
> 
> this by turning off and and on the laser stage control. I suspect that
> the 
> column may not have moved at all during the write, i.e. only one
> writefield 
> was written over and over again.
> John Paul, have you developed your write? Did it look like (2) hapened
> for 
> you also? Charis.
> 
> 
> and james:
> Greetings:
> The beam current values (BEAM I) have been falling consistently as the 10
> 
> and 30 um apertures are becoming blocked with hydrocarbons during the 
> writes. This is normal and is to be expected especially considering the 
> usage we have on the system.
> Currently the beam current for the 10 um aperture is 50% of when the 
> aperture disk was new.
> Currently the beam current for the 30 um aperture is 60 % of when the 
> aperture disk was new.
> 50 % of new maximum is the breakpoint to trigger the PM request. As
> planned 
> and previously announced, we will be doing a column clean with aperture 
> replacement PM including a NEW FE-gun tip replacement this next weekend.
> We 
> will be down Friday 10 AM through late Tuesday afternoon. No access no
> use 
> during this period. 40 hour bake out has been requested and will be done
> 
> concurrent with the Monday Holiday.
> The system will be requalified for performance and imaging resolution 
> before releasing it back to users. Please schedule your work
> accordingly.
> The report of the stage freeze is the first reported in some time... we 
> will monitor the situation. If you observe this give time for all macro
> and 
> other operations to be completed then attempt to move the stage a mm or 
> more using the Stage control. If this still does work press the STOP SIGN
> 
> in the Stage control window and it will stop all pending operations and 
> release the XYZ axis's.
> Back to green light mode... writes performed last night by dwshin came ot
> 
> very nice.
> Thank you,
> JWC
> 
> and hatice:
> For 30 micron aperture from 11:00 am to 12:30, current changed from from
> 
> 56pA to 63, 68, 113, 128, 110. It does not stabilize, it inreases and 
> decreases randomly.
> 
> 





More information about the raith mailing list